OpenStudy (anonymous):

Evan spent 20 hours doing homework last week. This week he spent 25 hours doing homework. He says that he spent 125% more time doing homework this week. Is he correct? Show your work to justify your decision.

OpenStudy (cowgirl1):

what do you think?

OpenStudy (study312):

what module is this? and is this 7th or 6th because I just did this but I can't remember what number it was

OpenStudy (cowgirl1):

its not correct, but what do you think it should be?

OpenStudy (study312):

what module is this? and is this 7th or 6th because I just did this but I can't remember what number it was. @ghostrider02

OpenStudy (michele_laino):

hint: we can write this computation: \[\Large 20 \times \frac{{125}}{{100}} = 25\] what can you conclude, please?

OpenStudy (anonymous):

7TH GRADE 4.05 and it equals 25

OpenStudy (michele_laino):

hint: if the \(125 \%\) of \(20\) is \(25\), then \((\)\(20+ 125\% \) of \(20\)\()\) is greater than \(25\)

OpenStudy (study312):

25 - 20 = ? can you tell me

OpenStudy (study312):

hello? still there?

OpenStudy (study312):

@ghostrider02?

OpenStudy (anonymous):

sorry I was outside it's 5

OpenStudy (jdoe0001):

so... is Evan correct?

OpenStudy (study312):

no

OpenStudy (jdoe0001):

well, I was wondering on @ghostrider02 's answer, since he's the one who asked

OpenStudy (study312):

25 - 20 = 5. 5 ÷ 20 = 0.25. 0.25 • 100 = 25. Evan worked 25% more not 125%